# Automated glenoid bone loss measurement and segmentation in CT scans for pre-operative planning in shoulder instability

### v1

@@ Abstract
Background: Reliable measurement of glenoid bone loss is essential for operative planning in shoulder instability, but current manual and semi-automated methods are time-consuming and often subject to inter-reader variability.
Purpose/Hypothesis: To develop and validate a fully automated, deep-learning pipeline for measuring glenoid bone loss on 3D computed tomography (CT) scans using linear-based, en-face view, and best-circle method. We hypothesized that automated measurements would perform comparable to expert surgeons.
Study Design: Cohort study (diagnosis); Level of evidence, 2.
Methods: Shoulder CT images of 91 patients (average age, 40 years; range, 14–89 years; 65 men) were retrospectively collected between January 2013 and March 2023, along with manually labeled glenoid and humerus segmentations, glenoid rim points, and percentage bone-loss measurements. Our fully automated multi-stage algorithm consists of three main stages: (1) Segmentation, where we developed a U-Net to automatically segment the glenoid and humerus; (2) anatomical landmark detection, where a second network predicts glenoid rim points; and (3) geometric fitting, where we applied a principal component analysis (PCA), projection, and circle fitting to compute the percentage of bone loss.
Results: The automated measurements showed strong agreement with consensus readings, exceeding surgeon-to-surgeon consistency (intraclass correlation coefficient [ICC] 0.84 vs 0.78). ICC values were 0.71 vs 0.63 for the low-bone-loss subgroup and 0.83 vs 0.21 for the high-bone-loss group ($P<.001$). For the classification task of assigning each patient to low-, medium-, or high–bone-loss severity subgroups, the pipeline’s recall (sensitivity) was 0.714 (5/7) for the low-severity group and 0.857 (6/7) for the high-severity group, with no instances of misclassifying low as high or vice versa.
Conclusion: We developed a novel, time-efficient, fully automated, deep learning–based pipeline for glenoid bone-loss measurement on 3D volume-rendered CT. The pipeline shows strong agreement with surgeon measurements, good reliability, and high recall for the low- and high-bone-loss subgroups. Therefore, it may serve as a clinically reliable tool to assist clinicians with preoperative planning for shoulder instability, particularly for screening patients with high bone loss, while still requiring human review for cases of moderate bone loss. We are releasing our model and dataset at https://github.com/Edenliu1/Auto-Glenoid-Measurement-DL-Pipeline.
Keywords: glenoid bone loss; deep learning; automated measurement; CT; segmentation; landmark
1 Background
@@
Many ideas have been proposed to solve these weaknesses; only a few are fully automated. Haimi’s method segmented the glenoid using a Hounsfield units (HU) threshold within an interval, applied PCA to reorient the glenoid slices to an en-face view, and fit a circle using a few predetermined clock positions on the glenoid segmentation [Haimi2024Automated_2d]. The flaws of this method are that HU-based segmentation may inaccurately include pixels of arthritis that fall within the interval; PCA is not the optimal solution for finding the en-face view because it can be largely influenced by cupping bone loss, which is discussed later in the paper; and the bone loss measurement was computed over a selected 2D slice of a CT scan, which many studies have found can underestimate the bone loss[Huijsmans2011]. Zhao et al.’s method requires a healthy contralateral side for the bone loss computation, while many CT scans include only one side[Zhao2023].
@@
We propose a fully automated, deep learning–based pipeline for 3D CT glenoid segmentation and measurement of glenoid bone loss using a linear-based, en-face view and the best-fit circle method. The glenoid bone is obtained using a deep learning model that is not subect to shoulder arthritis. A better approach is used to find the en face view, relying only on information from the intact side of the glenoid rim. Bone loss is measured on a 3D bone-rendered image rather than a 2D slice to avoid underestimation. A healthy contralateral side is not needed for this pipeline. Moreover, we assessed agreement using intraclass correlation coefficients (ICCs) to ensure that measurements to quantify inter-rater agreement across raters (algorithm vs. gold standard), which matters for understanding reliability of algorihtm for making surgical decisions. We also analyzed three bone-loss severity subgroups to approximate surgical decision management. We found that our pipeline for automating glenoid bone loss achieves higher agreement and correlation than the human baseline and demonstrates good reliability, especially in subgroups with low and high severity of bone loss. It is clinically reliable for assisting physicians in preoperative planning.
@@
The purpose of this study is to develop and validate a fully automated, deep-learning pipeline for measuring glenoid bone loss on 3D scans using glenoid rim points, geometric fitting, and a principle component analysis, projection, and circle fitting to compute the percentage of bone loss. We hypothesize that automated measurements would demonstrate comparable performance to surgeon measurements.
@@
Figure 1: Pipeline for automated glenoid bone loss measurement. (1) The segmentation model delineates the glenoid articular surface; (2) Rim-UNet reads the segmentation mask and predicts rim points; (3) singular value decomposition (SVD) estimates the rim-plane normal, defining the en face view; (4) the segmentation mask and rim points are projected onto the en-face plane; and (5) a 2D circle is fitted to the rim points. Bone loss is computed as $100\times B/A$, where $B$ is the glenoid-defect length and $A$ is the fitted-circle diameter.
@@ 2.1 Patient Data
Our dataset of 3D shoulder CTs were retrospectively collected from patients between January 2013 and March 2023. The training dataset was sampled from this entire dataset shoulder CT dataset and was screened and labeled by an orthopedic surgeon (over 15 years of experience). The test dataset was sampled from a cohort within the same dataset but comprising healthy patients and patients who had undergone shoulder stabilization surgery, and it was screened and labeled by two experienced surgeons. Exclusion criteria included: (1) patients with posterior glenoid bone loss, (2) patients with a shoulder implant, and (3) patients with severe arthritis (measurement not feasible for physicians).
@@
We stratified the screened dataset and randomly divide it into a 20% training data set and a 80% test data set, resulting in 21 patient CTs in the test data set and 77 in the train data set, as shown in Table 1. Demographic characteristics and scanne’s information of the trainning and testing cohorts can be found in Table 2.
@@
Characteristic Train (n = 61) Test (n = 21) Total (n = 92)
@@
SOMATOM Force 18 (30%) 9 (43%) 31 (34%)
Discovery CT750 HD 15 (25%) 10 (48%) 26 (28%)
SOMATOM Definition Flash 10 (16%) 2 (10%) 16 (17%)
Revolution HD 10 (16%) 0 (0%) 11 (12%)
SOMATOM Definition AS 4 (7%) 0 (0%) 4 (4%)
LightSpeed VCT 3 (5%) 0 (0%) 3 (3%)
@@ 2.2 Automated Glenoid Bone Loss Measurement Pipeline
We introduce a fully automated, end-to-end deep learning-based pipeline for segmenting the glenoid articular surface and measuring bone loss on 3D reconstructed CT using linear-based, en-face view, and best-fit circle method. The pipeline follows the same procedure applied by the clinician, as described in the introduction. Below is a brief pipeline description: 1. The trained deep learning segmentation model reads the shoulder CT scan and outputs the left/right glenoid and humerus bone loss labels. 2. The glenoid bone label is flipped along the depth dimension if detected as a Right glenoid. 3. The trained Rim-UNet reads the preprocessed binary segmentation map from the segmentation model and predicts the rim point sets. The predicted rim points are then back-registered to the original segmentation map 4. Given 3-D rim-skeleton points, we estimate a best-fit plane via SVD/PCA on the point cloud, taking the direction normal to the top two largest variance directions as the en-face view. 5. Glenoid segmentation and rim points are projected onto this plane. A 2-D geometric circle is fit by minimizing the radial least-squares objective $J(c,r)=\sum_{i}(\lVert p_{i}-c\rVert_{2}-r)^{2}$ using BFGS (c = center of circle coordinate; r = radius of the circle; p = rim point sets). Based on the clinical suggestion of the diameter of the fitted circle usually being 2/3 of the glenoid height, we did an optimization test using a grid search around the hyperparameter of 2/3 on the train set. Here we explicitly set the radius to be 0.6955/2 based on the analysis (Fig. 3). 6. Bone loss is measured by dividing the glenoid defect (B) by the diameter of the fitted circle (A): $(B/A)\times 100\%$. Following Hami’s method, determining B, many radial lines are drawn from the glenoid defect to the anterior margin of the circle, and the longest line represents the length of the glenoid defect [Haimi2024Automated_2d]. A general workflow of this fully automated pipeline is outlined in the flowchart provided in Figure 1.
@@ 2.3 Data Annotation
Segmentation Annotation: The annotation of the scapulae and humerus bones for the entire dataset was completed under the guidance of an orthopedic surgeon with more than 10 years of experience in sports medicine. We segmented the thin layer at the top of the glenoid fossa, which experts would consider a part to measure bone loss in the glenoid after finding the en-face view. All segmentation annotations are refined and approved by one surgeon for the training dataset and two for the test dataset.
Landmark annotation: A varied number of landmarks are selected on the surface of the posterior inferior rim on the annotated binary segmentation mask of the glenoid bone in 3D slicer, which will be used for training the model for circle fitting. The rim curve doctor refers to for circle fitting is shorter for patients with a lot of glenoid bone loss, but longer for patients with little bone loss. Therefore, we need to train a model that can output data accommodating different amounts of bone loss, which can output different sets of rim points. The instruction for selecting landmarks that best represent the intact posterior/inferior glenoid rim is given to researchers and surgeons.
@@
    Rotate the glenoid surface mask until the en-face view is normal to the labeler [Sugaya2003]
@@
    Select a landmark from what is considered 3 o’clock based on the labeler’s experience that coincided with the intact posterior-inferior rim [Sugaya2003]
@@
    Following the outermost rim, select the landmark until the curve meets the inferior defect aspect [Chen2022]
Similar to segmentation annotation, we manually label the landmark, which is then refined and approved by an experienced surgeon. One surgeon refined and approved the train dataset, while two surgeons refined and approved the test dataset.
2.4 Ground-Truth Generation
The ground truth for the segmentation model does not require any pre-processing, as the foreground binary mask can be used directly for training. In contrast, the ground truth for the rim model requires pre-processing, as there isn’t a fixed number and location of landmarks that can represent the curve for circle fitting across different level of bone loss. To enable the model to learn from the selected landmarks and generate information that can be easily post-processed for circle fitting, we follow these steps: 1. For consistency, the landmarks for each case were resampled to 30 landmarks on the 3D slicer. 2. Using the Markups to Model module in 3D slicer, the landmark markup is converted to a tube with a radius of 1.00 mm with a curve-type cardinal spline 3. The rim tube is then skeletonized into a thin curve with a thickness of 1 pixel. An isotropic 3D Gaussian kernel $\sigma=1\,\mathrm{mm}$ is applied to the thin skeletonized curve to regress a heatmap in the range of [0,1].
@@ 2.5 Model’s Architecture
Segmentation backbone (nnU-Net / TotalSegmentator). We only have 61 volumes in training data, which is considered a low number of images for deep learning training. To address this data scarcity, we conducted a two-stage fine-tuning of the public TotalSegmentator model, built on nnU-Net [Isensee2020] [Wasserthal2023]. Transfer learning is a highly data-efficient approach for training models in small subsets of data [Pan2010], and TotalSegmentator is a multi-label segmentation model trained on 104 anatomic structures across 1204 CT examinations, encompassing the scapula and humerus bones. We believe that fine-tuning Totalsegmentator can transfer its rich anatomical knowledge of these bones in the CT domain to our task of segmenting four target labels: left/right glenoids and left/right humerus bones.
In the second stage, we further fine-tuned the model on our subdata set to capture finer substructures, as described on the glenoid articular surface[Chlebus2018]. We collected a smaller dataset under the guidance of an orthopedic surgeon, comprising 98 CT scans and their corresponding left/ right glenoid articular surfaces and humerus bone loss, as shown in Table 1. The training plan used to fine-tune the first-stage model is identical to the one used in the first stage. We used the same loss function for both stages, L = LCE + LDICE, which combines entropy loss and the dice loss. We will release the exact plan/config JSON, trained weights, and scripts for full reproducibility (GitHub link coming up).
@@
Rim-UNet: Heatmap regression model for circle fitting. Manual measurement of glenoid bone loss in 3D CT scans by doctors in general requires (1) finding an en-face view, (2) projecting the 3D image in that direction, and (3) drawing a best-fit circle to the outermost rim/contour [Sugaya2003]. This pipeline would require solving two values: the direction vector at which a doctor would think makes the glenoid surface normal to their eye(en-face view), and the rim curve at the intact glenoid contour (excluding bone loss areas), which they refer to for circle fitting. Inspired by the Constellation Technique (CST), which labels the glenoid rim regardless of en face orientation, we introduce a novel approach to train a deep learning model for the best-fit circle on the glenoid contour, solving both tasks above with a single model and a small dataset. Rim data for training is selected first, reorienting the en-face view, and then, on that plane, select the outermost rim, encompassing information about both rim and the en-face view coordinates. Therefore, RimNet can skip the en-face view localizing and directly predict the outermost rim contour that the doctors use for circle fitting in 3D space. The en-face view direction vector can then be calculated with SVD on the curve points.
The RimNet is adapted from the structure of a vanilla U-net tailored for the heat-map regression thin-curve task. In addition to a basic encoder-decoder structure, we used strided convolution $(2,2,1)$ to replace the max-pooling for downsampling in the first two stages to preserve the details of the plane for our anistropic medical images [Isensee2020]. Moreover, for each convolution, we use Instance Normalization for medical images in small batches [Verweij2020] and LeakyReLU activations with $\alpha=0.01$ to preserve all gradients flowing to the prediction of our sparse target output [Maas2013]. We used adaptive wing loss for our landmark localization task, as a greater penalty is needed for medium to small errors. [Feng2018Wing][Wang2019AdaptiveWing].
[…]
2.6 Experimental Plan

### v2

# Fully Automated Deep Learning–Based Glenoid Bone Loss Measurement and Severity Stratification on 3D CT in Shoulder Instability

@@ Abstract
Purpose: To develop and validate a fully automated, deep-learning pipeline for measuring glenoid bone loss on 3D CT scans using linear-based, en-face view, and best-circle method.
Material and Methods: Shoulder CT scans of 81 patients were retrospectively collected between January 2013 and March 2023. Our algorithm consists of three main stages: (1) Segmentation, where we developed a U-Net to automatically segment the glenoid and humerus; (2) anatomical landmark detection, where a second network predicts glenoid rim points; and (3) geometric fitting, where we applied a principal component analysis (PCA), projection, and circle fitting to compute the percentage of bone loss. The performance of the pipeline was evaluated using DSC for segmentation and MAE and ICC for bone-loss measurement; intermediate outputs (rim point sets and en-face view) were also assessed.
Results: Automated measurements showed strong agreement with consensus readings, exceeding surgeon-to-surgeon consistency (ICC 0.84 vs 0.78 for all patients; ICC 0.71 vs 0.63 for low bone loss; ICC 0.83 vs 0.21 for high bone loss; $P<.001$). For the classification task of assigning each patient to different bone loss severity subgroups, the pipeline’s sensitivity was 71.4% for the low-severity group and 85.7% for the high-severity group, with no instances of misclassifying low as high or vice versa.
Conclusion: A fully automated, deep learning–based pipeline for glenoid bone-loss measurement on CT scans can be a clinically reliable tool to assist clinicians with preoperative planning for shoulder instability. We are releasing our model and dataset at https://github.com/Edenliu1/Auto-Glenoid-Measurement-DL-Pipeline.
Abbreviations: ICC = intraclass correlation coefficient, DSC = dice similarity coefficient, MAE = mean absolute Error
Summary: A fully automated, deep learning–based pipeline can accurately predict glenoid bone loss measurements and classify bone loss severity, providing clinical decision support to assist clinicians with preoperative planning for shoulder instability.
Key Points:
The pipeline showed good overall reliability relative to the physicians’ consensus and slightly outperformed the inter-reader (human–human) baseline (ICC 0.84 vs 0.78). The ICCs were 0.71 vs 0.63 for the low–bone-loss subgroup and 0.83 vs 0.21 for the high–bone-loss subgroup.
For classifying bone-loss severity, the model achieved a sensitivity of 71.4% in the low-bone-loss subgroup and 85.7% in the high–bone-loss subgroup, indicating robust performance in these two groups.
The algorithm’s strong classification performance in the low- and high-severity subgroups and its higher agreement with physician consensus than the inter-reader baseline suggest that the algorithm is clinically reliable for automating quantitative measurements, while human review remains important for intermediate-severity cases.
1 Introduction
@@
Many ideas have been proposed to solve these weaknesses; only a few are fully automated. Haimi’s method segmented the glenoid using a Hounsfield units (HU) threshold within an interval, applied PCA to reorient the glenoid slices to an en-face view, and fit a circle using a few predetermined clock positions on the glenoid segmentation [Haimi2024Automated_2d]. The flaws of this method are that HU-based segmentation may inaccurately include pixels of arthritis that fall within the interval; PCA is not the optimal solution for finding the en-face view because it can be largely influenced by cupping bone loss, which is discussed later in the paper in Appendix C; and the bone loss measurement was computed over a selected 2D slice of a CT scan, which many studies have found can underestimate the bone loss[Huijsmans2011]. Zhao et al.’s method requires a healthy contralateral side for the bone loss computation, while many CT scans include only one side[Zhao2023].
@@
We propose a fully automated, deep learning–based pipeline for 3D CT glenoid segmentation and measurement of glenoid bone loss using a linear-based, en-face view and the best-fit circle method. The glenoid bone is obtained using a deep learning model that is not subect to shoulder arthritis. A better approach is used to find the en face view, relying only on information from the intact side of the glenoid rim, which enables identification of the same en-face view of across different shapes of glenoid bone loss. Bone loss is measured on a 3D bone-rendered image rather than a 2D slice to avoid underestimation. A healthy contralateral side is not needed for this pipeline. Moreover, we assessed agreement using intraclass correlation coefficients (ICCs) to ensure that measurements to quantify inter-rater agreement across raters (algorithm vs. gold standard), which matters for understanding reliability of algorihtm for making surgical decisions. We also analyzed three bone-loss severity subgroups to approximate surgical decision management. We found that our pipeline for automating glenoid bone loss achieves higher agreement and correlation than the human baseline and demonstrates good reliability, especially in subgroups with low and high severity of bone loss. It is clinically reliable for assisting physicians in preoperative planning.
@@
The purpose of this study is to develop and validate a fully automated, deep-learning pipeline for measuring glenoid bone loss on 3D CT scans using glenoid rim points, geometric fitting, and a principle component analysis, projection, and circle fitting to compute the percentage of bone loss. We hypothesize that automated measurements would demonstrate comparable performance to surgeon measurements.
@@
Figure 1: Pipeline for automated glenoid bone loss measurement. (1) The segmentation model delineates the glenoid articular surface; (2) RimU-Net reads the segmentation mask and predicts rim points; (3) singular value decomposition (SVD) estimates the rim-plane normal, defining the en face view; (4) the segmentation mask and rim points are projected onto the en-face plane; and (5) a 2D circle is fitted to the rim points. Bone loss is computed as $100\times B/A$, where $B$ is the glenoid-defect length and $A$ is the fitted-circle diameter.
@@ 2.1 Patient Data
All CT scans were de-identified prior to analysis using a published DICOM metadata and burned-in text de-identification workflow [Macdonald2024]
Our dataset of 3D shoulder CTs were retrospectively collected from patients between January 2013 and March 2023. The training dataset was sampled from this entire dataset shoulder CT dataset and was screened and labeled by an orthopedic surgeon (over 15 years of experience). The test dataset was sampled from a cohort within the same dataset but comprising healthy patients and patients who had undergone shoulder stabilization surgery, and it was screened and labeled by two experienced surgeons. Exclusion criteria included: (1) patients with posterior glenoid bone loss, (2) patients with a shoulder implant, and (3) patients with severe arthritis (measurement not feasible for physicians). Further dataset details are provided in Appendix A.
@@
We stratified the screened dataset and randomly divide it into a 20% training data set and a 80% test data set, resulting in 21 patient CTs in the test data set and 60 in the train data set, as shown in Table 1. Demographic characteristics and scanner’s information of the trainning and testing cohorts can be found in Table 2.
@@
Characteristic Train (n = 60) Test (n = 21) Total (n = 81)
@@
SOMATOM Force 18 (30%) 9 (43%) 27 (33%)
Discovery CT750 HD 15 (25%) 10 (48%) 25 (31%)
SOMATOM Definition Flash 9 (15%) 2 (10%) 11 (14%)
Revolution HD 10 (17%) 0 (0%) 10 (12%)
SOMATOM Definition AS 4 (7%) 0 (0%) 4 (5%)
LightSpeed VCT 3 (5%) 0 (0%) 3 (4%)
@@ 2.2 Automated Glenoid Bone Loss Measurement Pipeline
We introduce a fully automated, end-to-end deep learning-based pipeline for segmenting the glenoid articular surface and measuring bone loss on 3D reconstructed CT using linear-based, en-face view, and best-fit circle method. The pipeline follows the same procedure applied by the clinician, as described in the introduction. Below is a brief pipeline description: 1. The trained deep learning segmentation model reads the shoulder CT scan and outputs the left/right glenoid and humerus bone loss labels. 2. The glenoid bone label is flipped along the depth dimension if detected as a Right glenoid. 3. The trained landmark neural network model (RimU-Net) reads the preprocessed binary segmentation map from the segmentation model and predicts the rim point sets. The predicted rim points are then back-registered to the original segmentation map 4. Given 3-D rim-skeleton points, we estimate a best-fit plane via SVD/PCA on the point cloud, taking the direction normal to the top two largest variance directions as the en-face view. 5. Glenoid segmentation and rim points are projected onto this plane. A 2-D geometric circle is fit by minimizing the radial least-squares objective $J(c,r)=\sum_{i}(\lVert p_{i}-c\rVert_{2}-r)^{2}$ using BFGS (c = center of circle coordinate; r = radius of the circle; p = rim point sets). Based on the clinical suggestion of the diameter of the fitted circle usually being 2/3 of the glenoid height, we did an optimization test using a grid search around the hyperparameter of 2/3 on the train set. Here we explicitly set the radius to be 0.6955/2 based on the analysis (Fig. 3). 6. Bone loss is measured by dividing the glenoid defect (B) by the diameter of the fitted circle (A): $(B/A)\times 100\%$. Following Hami’s method, determining B, many radial lines are drawn from the glenoid defect to the anterior margin of the circle, and the longest line represents the length of the glenoid defect [Haimi2024Automated_2d]. A general workflow of this fully automated pipeline is outlined in the flowchart provided in Figure 1.
@@ 2.3 Data Annotation
Segmentation annotation: The annotation of the scapulae and humerus bones for the entire dataset was completed under the guidance of an orthopedic surgeon with more than 10 years of experience in sports medicine. We segmented the thin layer at the top of the glenoid fossa, which experts would consider a part to measure bone loss in the glenoid after finding the en-face view. All segmentation annotations were refined and approved by one surgeon for the training dataset and by two surgeons for the test dataset.
Landmark annotation: A variable number of landmarks are selected on the posterior–inferior rim surface of the annotated binary glenoid segmentation mask in 3D Slicer, and these landmarks are used to train the circle-fitting model. The rim curve used for circle fitting is shorter in patients with substantial glenoid bone loss and longer in patients with minimal bone loss. Therefore, we train a model that can accommodate varying degrees of bone loss by outputting different sets of rim points. Researchers and surgeons are provided with instructions for selecting landmarks that best represent the intact posterior–inferior glenoid rim.
@@
    Rotate the glenoid surface mask until the en-face view is perpendicular to the labeler’s viewing plane [Sugaya2003].
@@
    Based on the labeler’s experience, select a landmark at the approximate 3 o’clock position that corresponds to the intact posterior–inferior rim [Sugaya2003].
@@
    Following the outermost rim, continue selecting landmarks until the curve reaches the inferior aspect of the defect [Chen2022].
Similar to the segmentation annotations, landmarks were manually labeled and then refined and approved by an experienced surgeon. One surgeon refined and approved the training dataset, whereas two surgeons refined and approved the test dataset.
@@ 2.4 Model’s Architecture
Segmentation backbone (nnU-Net / TotalSegmentator). We only have 61 volumes in training data, which is considered a low number of images for deep learning training. To address this data scarcity, we conducted a two-stage fine-tuning of the public TotalSegmentator model, built on nnU-Net [Isensee2020] [Wasserthal2023]. Transfer learning is a highly data-efficient approach for training models in small subsets of data [Pan2010], and TotalSegmentator is a multi-label segmentation model trained on 104 anatomic structures across 1204 CT examinations, encompassing the scapula and humerus bones. We believe that fine-tuning Totalsegmentator can transfer its rich anatomical knowledge of these bones in the CT domain to our task of segmenting four target labels: left/right glenoids and left/right humerus bones (Appendix A).
RimU-Net: Heatmap regression model for circle fitting.
RimU-Net is a U-Net based heatmap regression network designed for thin-curve localization. Given the glenoid surface segmentation, it predicts a heatmap/point set corresponding to the intact posterior–inferior rim, which is then used to estimate the best-fit circle. Architecture and training details are provided in Appendix B.
@@
Manual measurement of glenoid bone loss in 3D CT scans by doctors in general requires (1) finding an en-face view, (2) projecting the 3D image in that direction, and (3) drawing a best-fit circle to the outermost rim/contour [Sugaya2003]. This pipeline would require estimating two quantities: the direction vector at which a doctor would think makes the glenoid surface normal to their view (en-face view), and the rim curve at the intact glenoid contour (excluding bone loss areas), which they refer to for circle fitting. Inspired by the Constellation Technique (CST) [Chen2022], which labels the glenoid rim regardless of en-face orientation, we introduce a novel approach to train a deep learning model to estimate the best-fit circle on the glenoid contour, solving both tasks above with a single model using a small dataset. Rim data for training is selected first, reorienting the en-face view, and then, on that plane, select the outermost rim, encompassing information about both rim and the en-face view coordinates. Therefore, RimNet can skip the en-face view localizing and directly predict the outermost rim contour that the doctors use for circle fitting in 3D space. The en-face view direction vector can then be calculated with SVD on the curve points.
2.5 Evaluation